### v1

# Nested Apparent Horizons and Quantized Separation from Intense Hawking Backreaction

Abstract
When Hawking radiation from a rotating or non-rotating black hole becomes sufficiently intense, its own stress–energy can no longer be treated as a perturbation on a fixed background. In this regime the outgoing flux may generate an additional, transient trapping surface exterior to the original event horizon. Using a simple spherically symmetric semi-classical model we demonstrate that strong outgoing null energy can create nested apparent horizons, a feature reminiscent of the Penrose process but mediated by quantum backreaction. The effect is illustrated using a smooth Vaidya-type mass profile, and conditions for bifurcation and merger of horizons are derived. We further propose that the separation between nested horizons may obey a discrete quantization rule analogous to the Bohr–Sommerfeld condition, suggesting a geometric route toward quantum-gravity discreteness.
I Introduction
Hawking radiation [1, 2] provides a semi-classical bridge between general relativity and quantum field theory in curved spacetime. In the standard treatment, the stress tensor $\langle T_{\mu\nu}\rangle_{\text{ren}}$ of the emitted flux is assumed small enough that its gravitational backreaction on the geometry may be neglected. However, when the emission rate becomes extreme—for example near the endpoint of evaporation or in superradiant regimes close to the Kerr ergosphere—the outgoing energy density can locally rival the curvature scale, requiring a fully dynamical treatment.
In such situations the distinction between a global event horizon and local apparent or trapping horizons [3, 4] becomes essential. The aim of this work is to explore, within a minimal toy model, how strong outgoing radiation can generate a second apparent horizon exterior to the original one, leading to a transient “nested-horizon” configuration.
II Backreaction and the Vaidya framework
A convenient setting for dynamical spherically symmetric spacetimes with null flux is the Vaidya metric,
$ds^{2}=-\left(1-\frac{2M(u)}{r}\right)du^{2}-2\,du\,dr+r^{2}d\Omega^{2},$ (1)
where $u$ is retarded time and $M(u)$ is the Bondi mass observed at infinity. The associated stress tensor $T_{uu}=-(1/4\pi r^{2})\,dM/du$ describes outgoing null radiation. Local trapping surfaces are defined by the vanishing of the outgoing null expansion $\Theta_{\ell}=0$, yielding the apparent-horizon condition
$r_{\rm AH}(u)=2M_{\rm MS}(r,u),$ (2)
where $M_{\rm MS}$ is the Misner–Sharp mass inside radius $r$.
When the flux is gentle, $M_{\rm MS}(r,u)$ increases monotonically with $r$ and Eq. (2) has a single root. A sufficiently compact, energetic shell of outgoing radiation can, however, deform $M_{\rm MS}(r,u)$ so that Eq. (2) admits multiple roots, corresponding to nested apparent horizons.
III Smooth shell toy model
To illustrate this behavior, consider a continuous mass profile
$m(r,t)=M_{0}+M_{s}\,\frac{1+\tanh[(r-r_{s})/\delta]}{2},$ (3)
representing a black hole of mass $M_{0}$ surrounded by an outgoing null shell of energy $M_{s}$ centered at $r_{s}$ with thickness $\delta$. Apparent horizons satisfy
2m(r,t)=r.2\,m(r,t)=r. (4)
For small $M_{s}$ the single root $r=2M_{0}$ persists. As $M_{s}$ increases beyond a critical value $M_{s}^{\rm crit}$ (dependent on $\delta$ and $r_{s}$), Eq. (4) yields two or three intersections (Fig. 1), corresponding to an inner and an outer trapped surface. The outer root roughly follows $r_{\rm outer}\approx 2(M_{0}+M_{s})$, consistent with the heuristic step-function model.
Figure 1: Example of nested apparent horizons for $M_{0}=1$, $r_{s}=2.5$, $\delta=0.05$, and increasing shell energy $M_{s}$. The intersection of $2m(r)$ with $r$ gives the apparent horizons. For $M_{s}\lesssim 0.2$ there is a single root; for $M_{s}\gtrsim 0.3$ two additional roots appear, representing a transient outer horizon due to intense outgoing flux.
As the shell propagates outward, the outer apparent horizon expands and eventually vanishes once the shell radius exceeds the trapping condition, leaving only the inner horizon. In a fully time-dependent picture these surfaces merge and annihilate, in agreement with numerical studies of dynamical horizons [4, 5].
IV Quantized separation hypothesis and quantum-gravity implications
The nested–horizon configurations obtained above arise classically from non-linear backreaction between the Hawking flux and the background curvature. If one accepts that horizon area and surface gravity are adiabatic invariants, the possibility arises that the separation between successive apparent horizons may itself be subject to a discrete spectrum, analogous to quantized orbital radii in atomic physics.
IV.1 Bohr–Sommerfeld quantization of horizon motion
In semiclassical gravity, the product of surface gravity $\kappa$ and horizon area $A$ plays the role of an action variable. Following Bekenstein’s adiabatic–invariant argument [9], and subsequent formulations by Barvinsky and Kiefer [10] and Maggiore [11], one may postulate a Bohr–Sommerfeld–type condition
$\oint\kappa\,dA=2\pi n\hbar,\qquad n\in\mathbb{Z}^{+},$ (5)
which yields a discrete spectrum of allowed areas
$A_{n}=\epsilon\,n\,\ell_{\rm P}^{2},\qquad\epsilon\sim 8\pi,$ (6)
with $\ell_{\rm P}$ the Planck length. If a radiating black hole temporarily supports two apparent horizons $(r_{1},r_{2})$, their areas $A_{i}=4\pi r_{i}^{2}$ could then satisfy
$\Delta A=A_{2}-A_{1}=n\,\epsilon\,\ell_{\rm P}^{2}.$ (7)
Equation (7) imposes a quantum restriction on the possible radial separation $\Delta r=r_{2}-r_{1}$, yielding
$r_{2}^{2}-r_{1}^{2}=2n\,\ell_{\rm P}^{2}.$ (8)
Thus the nested horizons behave as discrete “gravitational shells,” with quantized spacing analogous to the electronic orbitals of an atom.
IV.2 Effective potential and resonant states
Between $r_{1}$ and $r_{2}$, null geodesics experience an effective potential $V_{\rm eff}(r)$ associated with the local Misner–Sharp mass function. A semiclassical WKB quantization condition,
$\int_{r_{1}}^{r_{2}}\!\!\sqrt{E^{2}-V_{\rm eff}(r)}\,dr=n\pi\hbar,$ (9)
would then determine the discrete, metastable configurations where outgoing Hawking flux and gravitational confinement balance. The inner and outer horizons form a self-consistent “quantum cavity” for vacuum modes, whose resonant states could manifest as discrete energy or area levels. In this sense, the horizon pair acts as a gravitational analog of the atomic shell structure, and evaporation corresponds to transitions between these quantized levels.
IV.3 Relation to quantum-gravity proposals
This interpretation connects naturally with several strands of quantum-gravity research. Loop-quantum-gravity treatments yield a discrete area spectrum $A_{j}=8\pi\gamma\ell_{\rm P}^{2}\sqrt{j(j+1)}$ for half-integer spin labels $j$ [12]. String-inspired and holographic models also predict quantized horizon microstates with equally spaced entropy spectra. The nested-horizon picture may offer a geometric manifestation of these discrete degrees of freedom in a semiclassical spacetime. If horizon separations are indeed quantized, the smallest possible $\Delta r$ would be of order the Planck length, establishing a minimal “horizon spacing” and suggesting that the evaporation endpoint corresponds to a finite, Planck-scale remnant.
IV.4 Prospects
Detecting such quantization directly is beyond current experimental reach, but analogous features might appear in numerical simulations of quantum collapse or in the discrete spectra of quasinormal modes. The nested-horizon framework therefore provides a potential bridge between classical backreaction dynamics and quantum discreteness, offering an avenue toward identifying the geometric origin of Planck-scale quantization and the microscopic basis of black-hole entropy.
V Discussion
The presence of transient nested apparent horizons does not imply two event horizons: the event horizon is a global construct depending on the entire future of the spacetime. Nevertheless, such transient trapped regions could influence the local thermodynamics and entropy bookkeeping of highly radiative black holes, and may provide a bridge toward models of regular or bouncing geometries [6, 7, 8].
In the extreme regime where the Hawking flux itself sources significant curvature, semiclassical consistency may fail and a quantum-gravitational treatment will be necessary. Nonetheless, the present toy model captures the essential qualitative mechanism by which strong backreaction can momentarily cloak an evaporating black hole within a second, dynamically generated apparent horizon.
Acknowledgments
The author thanks the UCLA Samueli Engineering MS program for support and colleagues for helpful discussions on dynamical horizons and quantum backreaction.